OpenStudy (anonymous):

please help solve this question (show your work)!!!! use limit of x->a f'(a)=(f[x]-f[a])/(x-a) to find the derivative of f[x]=x^2 +4, a=1

OpenStudy (anonymous):

help!!!! please!!!!!

OpenStudy (anonymous):

Solve this: \[ \lim_{x \rightarrow a} \frac{f(x) - f(a)}{x - a} = \lim_{x \rightarrow 1} \frac{x^2+4 - (1^2+4)}{x - 1} = \lim_{x \rightarrow 1} \frac{x^2-1}{x-1}\]

OpenStudy (anonymous):